If ABC is a right isosceles triangle whose hypotenuse is 4 units long, how long is each of the legs?
Answer by John10(297)

If ABC is a right isosceles triangle whose hypotenuse is 4 units long, how long is each of the legs?
You know that the right isosceles triangle has one right triangle and two EQUAL legs.
Apply the Pythagorean Theorem of the right triangle
Let one leg is x and other leg is y
we know that x = y
so we will have 4^2 = x^2 + y^2
x^2 + y^2 = 16
x^2 + x^2 = 16 because x = y
2x^2 = 16
x^2 =8
x = sqrt(8)
so each leg will be sqrt(8) units
John10:)